# Tensor Network Computations That Capture Strict Variationality, Volume Law Behavior, and the Efficient Representation of Neural Network States

### v1

@@ Introduction.
Figure 4: (a) (1+1)D TN representation of the Floquet dynamics of an $L$-site kicked Ising chain, starting from a product state. TNF contraction direction is either along the spatial direction or inverse time directon (see inverse direction results in SM). Conventional MPS-MPO contraction is along the time direction. (b) and (c) show the dynamics of the half-system bipartite von Neumann entropy $S(t)$ of a 14-site chain at $(J,h,g)=(\pi/4,\pi/4,0.5)$ and $(0.7,0.5,0.5)$, respectively. Insets of (b) and (c) present the entanglement spectrum for the 40 largest eigenvalues of the half-chain reduced density matrix at $t=15$.
@@
To illustrate this numerically, we consider the Floquet dynamics of an $L$-site 1D kicked Ising chain [46, 47, 48]. Expressing the Floquet evolution operator $F$ as a matrix product operator (MPO) (see SM for details), the time evolution corresponds to a (1+1)D tensor network, shown in Fig. S2(a). The tunable parameters in the model include the Ising interaction strength $J$, homogeneous transverse field $g$ and longitudinal field $h$. We focus on two representative cases for $L=14$ (i) the maximally chaotic case [49], $(J,g,h)=(\pi/4,\pi/4,0.5)$, which exhibits volume law physics, and (ii) the less chaotic case, $(J,g,h)=(0.7,0.5,0.5)$, which shows sub-volume-law physics at short times. For each time step $t$, we compute the bipartite von Neumann entropy $S(t)=-\mathrm{Tr}\left(\rho_{A}\log{\rho_{A}}\right)$ of the state $|\Psi(t)\rangle=F^{t}|0\rangle^{\otimes n}$, where $\rho_{A}$ is the reduced density matrix of the half-chain. (Computing $\rho_{A}$ is not an efficient operation, but we do so to illustrate the physics being captured). To define a useful TNF, we construct an contraction path that leads to configuration dependent isometries: we consider two natural ones, namely (i) contraction in the inverse time direction, and (ii) (transverse) contraction along the spatial direction. We then compare against the standard MPS-MPO compression to a bond dimension $\chi$. Results from (i), as well as from amplitude-independent MPO-MPO compression in the inverse time direction, are shown in the SM.
@@
The evolution of $S(t)$ is depicted in the main panel of Fig. S2(b) and (c), with exact results included. In the maximally chaotic regime (parameter set $(J,g,h)=(\pi/4,\pi/4,0.5)$), which might be thought of as closest to the random circuit intuition above, the entanglement entropy grows linearly from the beginning until half-width saturation. The required conventional MPS bond dimension to describe the full dynamics grows exponentially with time (here for a finite $L=14$ chain the maximum $\chi$ needed would be $1282^{L/2}=$ 128). However, volume law physics is already recovered using a TNF with $\chi=2$. The $\chi=8$ TNF accurately captures the features of the full evolution, including the maximum point in the entropy at $t=8$. Similar results are seen for the TNF defined in the inverse time-direction (see SM), although it is quantitatively less accurate.
@@
In the less chaotic regime $(J,g,h)=(0.7,0.5,0.5)$, the entanglement entropy initially grows sublinearly up until $t=7$, as shown in Fig. S2(c). The TNF for any $\chi$ captures the initial growth and more entanglement than the conventional MPS of the same bond dimension. However, at small $\chi$ the TNF entropy saturates at too small a value relative to the exact result (although the exact result also saturates below the maximum value, see SM [si]). Nonetheless, the entanglement spectrum (insets of Fig. S2) for the TNF, even for small $\chi$, has a long tail and captures the broad features of the spectrum, rather than exhibiting the sharp cutoff of a conventional truncated MPS.
@@
However, if one allows for general tensor network geometries, it is straightforward to construct a TNF (without using isometries) that represents a neural network quantum state without the restriction of an area law. We consider the explicit example of a feed forward neural network (FNN), a popular universal architecture with $k$-layers of neurons, where the outputs of each layer $\{y^{(k-1)}_{i}\}$ form the inputs $\{x^{(k)}_{i}\}$ to the next layer, and the $i$-th neuron (activation) function in layer $k$ is $f^{(k)}_{i}(\{x^{(k)}_{j}\};\theta_{i}^{(k)})$ where $\theta_{i}^{(k)}$ are the neuron parameters. Assuming that all $f$ are implemented without recursion, then the computational graph of a FNN is a directed acyclic graph. Then we observe that (i) if the FNN is efficient, the classical binary circuit implementation of the neural network function is of polynomial size in logic and copy gates, where a binary logic gate is a tensor with two binary input legs and one binary output leg, while copy takes one binary input and returns two binary outputs (nonlinearity of the activation functions enters through the copy operator) (ii) given the input product state $|\vec{n}\rangle$ (i.e. a bitstring), the application of each gate $G$ produces another product state $|\vec{n}^{\prime}\rangle$ (i.e. another bitstring). Thus contracting the resulting tensor network in the appropriate order from inputs to the outputs (for example, using sparse tensor contraction, or by introducing SVDs to discover the product state structure of $|\vec{n^{\prime}}\rangle$, see SM for more details) can always be done efficiently.
@@ Acknowledgments.
In the Fig.4(c) in the main article where the dynamics is less chaotic at $(J,g,h)=(0.7,0.5,0.5)$, the half-chain entanglement entropy of the tensor network function appears to saturate at Floquet time $t=15$. To show that the exact entanglement entropy grows very slowly at long times, we present more calculation results up to Floquet time $t=90$ for $L=10$ and $L=14$ chains, as shown in Fig. S3.
@@
One can see that with the same $\chi$, the tensor network function contracted in the inverse direction captures more entanglement than the conventional MPS, as well as the Heisenberg-style MPS from MPO-MPO contraction. However, compared to the transverse contraction, inverse-time contraction has a lower accuracy at the early steps of the Floquet dynamics, showing a convergence in the entanglement entropy ahead of the true saturation for the maximally chaotic kicked Ising dynamics.
@@
In our first construction of the tensor network function representation of a feed forward neural network we use classical binary arithmetic using logic and copy gates. Here we describe in more detail this binary arithmetic and the building blocks (in terms of the tensors) of polynomial functions. (We can use polynomial functions to efficiently represent activation functions in a neural network under mild conditions). Note that we have chosen the construction below for simplicity and this is not the only binary logic construction of arithmetic: for more examples, we refer to [56, 57, 58, 59].

### v2

@@ Introduction.
Refer to captionRefer to caption
Figure 4: (a) (1+1)D TN representation of the Floquet dynamics of an $L$-site kicked Ising chain, starting from a product state. TNF contraction direction is either along the spatial direction or inverse time directon (see inverse direction results in SM [46]). Conventional MPS-MPO contraction is along the time direction. (b) and (c) show the dynamics of the half-system bipartite von Neumann entropy $S(t)$ of a 14-site chain at $(J,h,g)=(\pi/4,\pi/4,0.5)$ and $(0.7,0.5,0.5)$, respectively. Insets of (b) and (c) present the entanglement spectrum for the 40 largest eigenvalues of the half-chain reduced density matrix at $t=15$.
@@
To illustrate this numerically, we consider the Floquet dynamics of an $L$-site 1D kicked Ising chain [47, 48, 49]. Expressing the Floquet evolution operator $F$ as a matrix product operator (MPO) (see SM for details), the time evolution corresponds to a (1+1)D tensor network, shown in Fig. 4(a). The tunable parameters in the model include the Ising interaction strength $J$, homogeneous transverse field $g$ and longitudinal field $h$. We focus on two representative cases for $L=14$ (i) the maximally chaotic case [50], $(J,g,h)=(\pi/4,\pi/4,0.5)$, which exhibits volume law physics, and (ii) the less chaotic case, $(J,g,h)=(0.7,0.5,0.5)$, which shows sub-volume-law physics at short times. For each time step $t$, we compute the bipartite von Neumann entropy $S(t)=-\mathrm{Tr}\left(\rho_{A}\log{\rho_{A}}\right)$ of the state $|\Psi(t)\rangle=F^{t}|0\rangle^{\otimes L}$, where $\rho_{A}$ is the reduced density matrix of the half-chain. (Computing $\rho_{A}$ is not an efficient operation, but we do so to illustrate the physics being captured). To define a useful TNF, we construct an contraction path that leads to configuration dependent isometries: we consider two natural ones, namely (i) contraction in the inverse time direction, and (ii) (transverse) contraction along the spatial direction. We then compare against the standard MPS-MPO compression to a bond dimension $\chi$. Results from (i), as well as from amplitude-independent MPO-MPO compression in the inverse time direction, are shown in the SM [46].
@@
The evolution of $S(t)$ is depicted in the main panel of Fig. 4(b) and (c), with exact results included. In the maximally chaotic regime (parameter set $(J,g,h)=(\pi/4,\pi/4,0.5)$), which might be thought of as closest to the random circuit intuition above, the entanglement entropy grows linearly from the beginning until half-width saturation. The required conventional MPS bond dimension to describe the full dynamics grows exponentially with time (here for a finite $L=14$ chain the maximum $\chi$ needed would be $1282^{L/2}=$ 128). However, volume law physics is already recovered using a TNF with $\chi=2$. The $\chi=8$ TNF accurately captures the features of the full evolution, including the maximum point in the entropy at $t=8$. Similar results are seen for the TNF defined in the inverse time-direction (see SM [46]), although it is quantitatively less accurate.
@@
In the less chaotic regime $(J,g,h)=(0.7,0.5,0.5)$, the entanglement entropy initially grows sublinearly up until $t=7$, as shown in Fig. 4(c). The TNF for any $\chi$ captures the initial growth and more entanglement than the conventional MPS of the same bond dimension. However, at small $\chi$ the TNF entropy saturates at too small a value relative to the exact result (although the exact result also saturates below the maximum value, see SM). Nonetheless, the entanglement spectrum (insets of Fig. 4) for the TNF, even for small $\chi$, has a long tail and captures the broad features of the spectrum, rather than exhibiting the sharp cutoff of a conventional truncated MPS.
@@
However, if one allows for general tensor network geometries, it is straightforward to construct a TNF (without using isometries) that represents a neural network quantum state without the restriction of an area law. We consider the explicit example of a feed forward neural network (FNN), a popular universal architecture with $k$-layers of neurons, where the outputs of each layer $\{y^{(k-1)}_{i}\}$ form the inputs $\{x^{(k)}_{i}\}$ to the next layer, and the $i$-th neuron (activation) function in layer $k$ is $f^{(k)}_{i}(\{x^{(k)}_{j}\};\theta_{i}^{(k)})$ where $\theta_{i}^{(k)}$ are the neuron parameters. Assuming that all $f$ are implemented without recursion, then the computational graph of a FNN is a directed acyclic graph. Then we observe that (i) if the FNN is efficient, the classical binary circuit implementation of the neural network function is of polynomial size in logic and copy gates, where a binary logic gate is a tensor with two binary input legs and one binary output leg, while copy takes one binary input and returns two binary outputs (nonlinearity of the activation functions enters through the copy operator) (ii) given the input product state $|\vec{n}\rangle$ (i.e. a bitstring), the application of each gate $G$ produces another product state $|\vec{n}^{\prime}\rangle$ (i.e. another bitstring). Thus contracting the resulting tensor network in the appropriate order from inputs to the outputs (for example, using sparse tensor contraction, or by introducing SVDs to discover the product state structure of $|\vec{n^{\prime}}\rangle$, see SM for more details [46]) can always be done efficiently.
@@ Acknowledgments.
In the Fig. 4(c) in the main article where the dynamics is less chaotic at $(J,g,h)=(0.7,0.5,0.5)$, the half-chain entanglement entropy of the tensor network function appears to saturate at Floquet time $t=15$. To show that the entanglement entropy grows very slowly at long times, we present more calculation results up to Floquet time $t=90$ for $L=10$ and $L=14$ chains, as shown in Fig. S3.
@@
One can see that with the same $\chi$, the tensor network function contracted in the inverse direction captures more entanglement than the conventional MPS, as well as the Heisenberg-style MPS from MPO-MPO contraction. However, compared to the transverse contraction, inverse-time contraction has a lower accuracy at the early steps of the Floquet dynamics, showing a convergence of the entanglement entropy ahead of the true saturation time for the maximally chaotic kicked Ising dynamics.
@@
In our first construction of the tensor network function representation of a feed forward neural network, we use classical binary arithmetic using logic and copy gates. Here we describe in more detail this binary arithmetic and the building blocks (in terms of the tensors) of polynomial functions. (We can use polynomial functions to efficiently represent activation functions in a neural network under mild conditions). Note that we have chosen the construction below for simplicity and this is not the only binary logic construction of arithmetic: for more examples, we refer to [57, 58, 59, 60].